my slope is 3 through the point 2,11 and i have to find the equation of the line in slop-intercept form how do i do that?
Answer by checkley71(8403)

Y=mX+b IS THE LINE FORMULA WHERE m=SLOPE & b= Y INTERCEPT.
NOW REPLACE THE X & Y TERMS WITH 2 & 11 AND THE m WITH 3 & SOLVE FOR b.
11=3*2+b
b=11-6
b=5 THIS IS THE Y INTERCEPT.
Y=3X+5 THIS IS THE LINE FORMULA WITH A SLOPE OF 3 THROUGH (2,11).